OpenStudy (animalover02):

I WILL METAL!!!!!!!! which ordered pair is a solution to the linear inequality? y=4x is less than or equal to -6 A. (-2 , 4) B. (1 , -2) C. (1 , 3) D. (2 , 3)

OpenStudy (zacharydoc):

a

563blackghost (563blackghost):

@zacharydoc you can not be giving the answers away, you must help the owlet so that they may understand the question and how to solve. This is in violation of the code of conduct. Please refrain from this in the future.

OpenStudy (animalover02):

was it the right answer though???

563blackghost (563blackghost):

Is it \(\large\bf{y-4x \le -6}\) ? Cause you put `y=4x`

OpenStudy (animalover02):

yes thats correct how you wrote it.

OpenStudy (animalover02):

is the answer still A? or no?

563blackghost (563blackghost):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zacharydoc a \(\color{#0cbb34}{\text{End of Quote}}\) This is not correct. \(\huge\bf{4-4(-2) = 12}\) `12 is not less than -6`

OpenStudy (will.h):

u can subsitute the values of the point in the inequality and if it sets the inequality correctly then it would be one of the solutions

OpenStudy (animalover02):

I think C orD is right i think B is wrong.

563blackghost (563blackghost):

Lets try B. \(\huge\bf{-2-4(1) \le -6}\)

OpenStudy (animalover02):

B is wrong

OpenStudy (animalover02):

right b is wrong?

563blackghost (563blackghost):

Well try to simplify it. How do you know for sure? \(\huge\bf{-2-4(1) = -6}\) \(\huge\bf{-6 \le -6}\) This is true since it is `less than or equal to`.

OpenStudy (error1603):

Can I help BB?

OpenStudy (animalover02):

wait so B is correct??

563blackghost (563blackghost):

Yes it is.

OpenStudy (animalover02):

ok thanks:)

563blackghost (563blackghost):

np :)